OpenStudy (anonymous):

Find the complex zeros of the polynomial function. Write f in factored form. f(x)=x^3-5x^2+16x-30 Use the complex zeros to write f in factored form.

OpenStudy (jdoe0001):

using the "rational root test", you'd find that 3, is a root of it, a real root so dividing it for (x-3), that is $$ \bf \frac{x^3-5x^2+16x-30}{x-3} \implies x^2-2x+10\\ \text{so you have then}\\ f(x)=(x-3)(x^2-2x+10) $$ to get the 2 complex roots, I gather I'd use the quadratic formula on the trinomial

OpenStudy (anonymous):

The example shows f(x)=(x-2)(x-4-5i)(x-4+5i) for what the examples answer looks like

OpenStudy (jdoe0001):

hmm well

OpenStudy (jdoe0001):

$$\bf \frac{x^3-5x^2+16x-30}{x-3} \implies x^2-2x+10\\ \text{so you have then}\\ f(x)=(x-3)(x^2-2x+10)\\ x= \cfrac{ 2 \pm \sqrt { 4-40}}{2} \implies x= \cfrac{ 2 \pm \sqrt {-36}}{2}\\ x= \cfrac{ 2 \pm 6i}{2} \implies 1\pm 3i $$

OpenStudy (jdoe0001):

\(\large f(x)=x^3-5x^2+16x-30 \implies \bf (x-3)(x-1-3i)(x-1+3i)\)

OpenStudy (jdoe0001):

woops, got truncated :/

OpenStudy (anonymous):

I wish there was away I could take a picture of the exampes and post them

OpenStudy (jdoe0001):

\(\large f(x)=x^3-5x^2+16x-30 \\ \implies \bf (x-3)(x-1-3i)(x-1+3i)\)

OpenStudy (jdoe0001):

just take a screenshot, in paste it in Paint, and save, and here click on [Attach File]

OpenStudy (anonymous):

Find the complex zeros of the polynomial function. Write f in factored form. f(x)=x^3+8

OpenStudy (anonymous):

I don't know how to take a screen shot on a mac

OpenStudy (jdoe0001):

well, is usually easy to post the question in the channel, so we can all see it and help, and revise each other

OpenStudy (jdoe0001):

taking screenshots on a mac.... I think is about the same

OpenStudy (anonymous):

ok. I did post the next question on this feed

OpenStudy (jdoe0001):

well, I guess on the main channel :) either way, this one is pretty simple

OpenStudy (jdoe0001):

$$ \bf a^3+b^3 = (a+b)(a^2-ab+b^2)\\ x^3+8 \implies x^3+2^3\\ \implies (x+2)(x^2-2x+2^2) $$

OpenStudy (jdoe0001):

so, there's right off, the real positive root, (x+2) the rest, you can again, use the quadratic formula to squeeze out the complex ones

OpenStudy (anonymous):

First, find the real zero of the function by taking a real cubic root. Then divide f(x) by the linear corresponding to the real zero. Finally, find the complex zero of the resulting quotient using the quadratic formula.

OpenStudy (jdoe0001):

well, right, the thing is to find a real cubic root, usually you'd need to use the \(\bf \text{rational root test}\)

OpenStudy (jdoe0001):

the one you just posted is quite simple because there's an identity already for that form \(\bf a^3+b^3 = (a+b)(a^2-ab+b^2)\)

OpenStudy (anonymous):

it's f(x)=x^3+8

OpenStudy (jdoe0001):

\(\bf \color{blue}{a^3+b^3 = (a+b)(a^2-ab+b^2)}\\ x^3+8 \implies x^3+2^3\\ \implies (x+2)(x^2-2x+2^2) \)

OpenStudy (anonymous):

that is what I typed in

OpenStudy (jdoe0001):

well, yes, that's a factored form, from the trinomial, you can get the complex ones by using the quadratic formula like for the previous one

OpenStudy (anonymous):

when I typed that answer in it said it was wrong then it told me First, find the real zero of the function by taking a real cubic root. Then divide f(x) by the linear corresponding to the real zero. Finally, find the complex zero of the resulting quotient using the quadratic formula.

OpenStudy (jdoe0001):

yes, because you haven't use the quadratic formula to get the complex ones yes

OpenStudy (jdoe0001):

yet rather

OpenStudy (anonymous):

I don't know how to get that answer.

OpenStudy (jdoe0001):

\((x+2)(x^2-2x+2^2)\) so this is one root, a real positive one => (x+2) so the other 2 will come out of \(\bf (x^2-2x+2^2)\) use the quadratic formula on that one, solve for "x"

OpenStudy (anonymous):

I have to use websites to help me answer problems and the site I use when I type this in won't give me an option to solve for x

OpenStudy (jdoe0001):

so, just use the quadratic formula, you pretty much have the values to plug right in

OpenStudy (anonymous):

i don't know how to do it at all that is why I have to ask for help

OpenStudy (jdoe0001):

$$ \text{let's use the quadratic formula on}\\ \large { (x^2-2x+2^2) \implies (\color{red}{1}x^2\color{blue}{-2}x+\color{green}{2}^2)\\ x= \cfrac{ - \color{blue}{b} \pm \sqrt { \color{blue}{b}^2 -4\color{red}{a}\color{green}{c}}}{2\color{red}{a}} } $$

OpenStudy (jdoe0001):

so who goes where?

OpenStudy (jdoe0001):

see, rather :/ see who goes where? just plug them in and you'd get the values